M328K 51125 Spring ’97 T Th 11:00 - 12:30 RLM 6.120

Instructor: C Friedman RLM 11.140 471-5161 office hrs: T Th 2:30-3:30pm or by appt.

Text and Syllabus: Elementary Number Theory, 2nd ed.,Underwood Dudley, W.H.Freeman and Co. Syllabus; most of the text-a few sections probably omitted.

Exams: We will have exams on 2/13 (class 10; Thurs.), 3/27 (class 20: Thurs.), 5/10 (Final exam; Sat. 9-12am ). All exam grades will be curved. MAKEUPS WILL BE GIVEN VERY RELUCTANTLY

Homework: Homework will be assigned each Tuesday in class and will be due the next Tuesday. A serious effort should be made on the exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, all homework grades together will be counted as equivalent to one exam grade.

Final Grades: In computing a final grade, I will count the three tests and one total homework grade equally. (So there will be four grades to average.)